OpenStudy (anonymous):

What is the slope of the line that contains the points (-7, 5) and (-2, 3)?

Mehek (mehek14):

use the same formula as before \(\dfrac{3 - 5}{-2 - (-7)} = \dfrac{-2}{5}\)

OpenStudy (anonymous):

thanks again! (: @Mehek14

Mehek (mehek14):

\(\ddot\smile\)